### v1

@@ 4.4  Individual referees
Since perceptions and consequences of colorism and bias may differ by gender and skin tone of the referees [2, 9, 22, 37, 66], it is also interesting to explore individual—gender- or skin-tone-related—referee bias/favoritism. The ideal way for isolating personal-level effects would be to have data that links each foul to a specific referee. While each game features a three-person officiating crew, unfortunately, we cannot observe which referee is responsible for each call. Therefore, we cannot precisely link individual referees to the exact fouls they call, which would be ideal for isolating personal-level effects.
We, nevertheless, follow [55], who acknowledged a similar caveat in their data, to explore individual referees’ associations to possible bias. For that, we use data on individual referees who appeared in at least 1% of all possible games in the 2007-2010 period. We chose this period since it is the only period that has been consistently linked to overcorrection in the previous analyses. More specifically, we follow the approach of [58] by replacing the $\text{CrewDistance}_{ig}$ variable (the average absolute difference in skin tone (L* value) between a player and each member of the three-person referee crew) in equation (2) with an interaction term between a dummy for a specific referee and the absolute skin tone distance of each referee to a specific player, replacing $\gamma_{g}$ controlling for attendance and team performance up until the specific game. We run this estimation separately for each referee to obtain coefficients indicating how much, on average, referee r deviates from the other referees.
Figure 6 shows the results for each referee, where a triangle (woman) or circle (man) represent an estimate of the number of extra fouls per 40 minutes an individual WNBA referee calls based on the skin tone of players relative to the referee, with the bars representing the 95% confidence interval around these estimates. Referees are ordered along the x-axis by their estimated effect, from the lowest to the highest impact on foul-calling based on skin tone. The color of the point estimates and error bars corresponds to the referee’s representative skin tone.
Figure 6:
We find that out of 39 referees, six show significant bias (positive effect) against dissimilar players, and seven show significant favoritism (negative effect). For all the full sample period and the sub-sample periods, we find no significant association between their and the players’ tone color and the foul calls (see Appendix Figures C.1, C.2 and C.3). Likewise, no clear pattern emerges when looking at foul calling between women (triangles) and men (circles) referees. [^24] Including a set of interaction terms regarding referee crew gender when estimating equation (2), we do not find evidence in our setting to suggest that referee gender influences the expression of colorism (results are available upon the request).
@@ 5  Discussion
In contrast to the results for the NBA by [55], we do not find evidence of racial discrimination in the WNBA before the widespread media coverage that was provoked by the publication of the [55] results. Remarkably, however, we observe that immediately following the increased public awareness (2007-2010), a player earns fewer fouls when facing more referees from the opposite race. While the estimates are not significant at conventional levels when using the FairFace algorithm, it is precise enough to become significant in the specifications using race categorizations from human raters, and importantly, when limiting the data to observations for which race is unambiguous to both human raters and FairFace, and for the continuous variable of skin tone.
Overall, this highlights the critical distinction between race and skin tone. In the binary classification by human raters and FairFace, race and skin tone are inherently entangled and each evaluation relies on skin tone but also cues such as names and/or facial features to make racial assessments. In contrast, the continuous measure isolates skin tone. When doing so, significant results emerge, supporting the notion that skin tone is an important part of the mechanism and may warrant a reexamination of past empirical studies on racial discrimination.
The results in the NBA for the period immediately after the extensive media coverage are the same as in the WNBA, that is, players receive significantly fewer foul calls from the crew that has more opposite race referees. However, while the direction of a potential “media attention”-effect in the NBA and the WNBA is the same, the crucial difference between both settings is that before media exposure, there was evidence of racial discrimination in the NBA, while there was no such evidence in the WNBA. These ex-ante differences translate into zero bias in the NBA in the immediate post-awareness period, but overcorrection in the WNBA. However, and consistent across specifications: it appears that the overcorrection wore off over time, returning to zero-bias levels in the later post-awareness period (2011-2014). [^25] The difference in results between the FairFace algorithm and human raters raises important questions about the broader implications of using state-of-the-art algorithms to make judgments traditionally reserved for humans. While we believe the analytical benefits, particularly in terms of scalability, replicability, and objectivity far outweigh the potential drawbacks in this context, we acknowledge that these technologies introduce new ethical and practical considerations. Although a full discussion is beyond the scope of this paper, we believe there is a pressing need for ongoing dialogue about how best to integrate such advancements responsibly and effectively into empirical research.
@@ 5.1  Potential mechanisms
Our analyses allow us to probe deeper into the mechanisms of overcorrection. First, using the continuous variable of the skin tone, we find evidence that this overcorrection is driven by out-group favoritism rather than by in-group discrimination. This is in line with [26], who found a general pattern of favoritism rather than discrimination. Second, it is important to understand whether this out-group favoritism is intentional or not. For that, we dive into literature on the economics of discrimination, whose three main models include taste-based discrimination [11], statistical discrimination [10, 53], and implicit discrimination [12].
The first type of models, taste-based discrimination, assumes that individuals have certain preferences against a certain group of individuals. However, this type is highly unlikely in a competitive market of sports referees [23]. [55] also cast doubts about this mechanism in the NBA. In addition, our results regarding individual referees in the overcorrection period show that there are only very few referees with a significant bias, supporting the notion that taste-based discrimination is the unlikely candidate to explain our results. The second type of models, statistical discrimination, assumes asymmetric information where race is taken as a signal. As such, in the context of foul calls in the WNBA, where there is no problem of asymmetric information, this mechanism can be ruled out as well.
Consistent with other sports-related studies on referee bias, implicit discrimination is the most likely explanation for overcorrection in our case [55, 27, 54, 25]. This alignment is supported by the fact that referee decisions in professional sports often involve high cognitive load (e.g., time pressure) and a degree of ambiguity, primary environmental preconditions for implicit bias [12]. Given that we find evidence of out-group favoritism, we adjust the term from implicit discrimination to implicit favoritism. Such favoritism is unintentional and outside of the individual’s awareness. It may arise due to the unconscious association of a social group with different attributes and social norms. Social norms are especially important in the WNBA because of its strong emphasis on DEI. This commitment is apparent in the league’s programs, communications, and general atmosphere and thus may affect unconscious decisions that are taken in the type of split-second, high-pressure situations that appear in basketball games. [^26] It is worth mentioning a former player Candice Wiggins who claimed that the environment around the league was one of the reasons she chose to retire early. She mentioned feeling like an outsider in the WNBA as a heterosexual woman, noting that the league’s strong culture made it difficult for those who didn’t conform [24].
@@ 5.2  Limitations
Despite the strengths of our approach, two limitations merit consideration. First, although we use a continuous, perceptually grounded measure of skin tone and validate racial classifications using both human raters and the FairFace algorithm, these methods depend on visual cues and are subject to noise from image quality, lighting, and subjective interpretation. While steps like HSV filtering and k-means clustering mitigate these issues, they cannot fully eliminate them. Second, our analysis relies on referee crew composition rather than individual-level foul call attribution. Thus, we cannot precisely link individual referees to the exact fouls they call, which would be ideal for isolating personal-level effects.
@@ 6  Conclusion
While increased awareness is supposed to reduce racial bias in organizations that have experienced racial discrimination, this paper aimed to investigate the effect of increased awareness in non-biased organizations regarding race and skin tone. For that, we used widespread media coverage of the racial bias of referees in the NBA to test whether it had an effect on the decision-making of referees in the WNBA, an organization that is known for its strong position on equal rights and fight against discrimination. Using different state-of-the-art machine learning algorithms to define race and skin tone, our empirical analyses found that, as expected, WNBA referees were not racially biased in their foul calls’ decisions before the media explosion of the NBA results. However, an increased awareness caused overcorrection, according to which players received significantly fewer foul calls when facing more referees from the opposite race. We also found that this overcorrection disappeared in the long term.
While the external validity of our results is an open question, due to the unique nature of the WNBA, our results may still have insight into the implementation of DEI initiatives. While we do not take a normative stance on such programs, rather, we emphasize that their impact depends heavily on context. In settings where baseline levels of bias are already low, well-intentioned interventions may unintentionally disrupt the already unbiased behaviors and create pressure to overcorrect. Finally, while a significant body of economic research has explored out-group discrimination, our results call for a more comprehensive understanding of environments that may create overcorrection. Controlled experiments could test the generalizability of this phenomenon across other settings, while theoretical work could elucidate mechanisms linking group identity, salience, and behavior. Future research should aim to integrate these insights into models of social preferences, further enriching our understanding of identity-driven decision-making in both individual and institutional contexts.
@@
  Table C.1 Racial bias among WNBA referees: FairFace classification .22
  Table C.2 Racial bias among WNBA referees: Human raters .22
  Table C.3 Colorism among WNBA referees .23
  Figure C.1 Distribution of colorism by referee .23
  Figure C.2 Distribution of colorism by referee (2004–06) .24
  Figure C.3 Distribution of colorism by referee (2011–14) .24
@@ Appendix B Methodology
As mentioned in the main text, our third approach follows the “Image-to-Data Pipeline” in [4], dividing the process into two main parts, i.e., face segmentation and skin tone classification. The detailed process is described here:
@@ B.1 Face segmentation
B.1.2 Isolating skin components [^27] We do, however, omit the Conditional Random Field (CRF), which is used by [4] to refine the skin mask by applying a CRF module, which predicts the labels of neighboring pixels (i.e., whether they are predicted to be skin or not skin) to produce a more fine-grained segmentation result. In our case, it is unnecessary since all our photos are human, and we will have additional skin mask refinement later (see feature processing). The added complexity a CRF module would bring is not a value add in our case scenario.: We use the Fully Connected Convolutional Neural Network (FC-CNN) method to extract a skin mask from faces. This involves applying a fully-convolutional neural network (FCN), a type of convolutional neural network (CNN) where the last fully connected layer is substituted with a convolutional layer that can capture the locations of the predicted labels. This enables the prediction of peripheral landmarks, such as the edges of the facial skin area, eyes, nose, and mouth.
@@
B.1.4 Creating face masks: Utilizing the predicted landmarks from the FC-CNN, we further isolate the targeted facial region by extracting a skin “mask” from these predicted landmarks using the built-in Convex Hull method from SciPy’s Python library. The Convex Hull method generates the hull, and the skimage’s draw polygon function creates a binary mask isolating the face region.
@@ B.2.1 Skin Extraction
B.2.1.6 Conversion to BGR Color Space: The extracted skin regions were finally converted back to the BGR color space for further analysis.
@@ Appendix C Results
Change in coefficient, -0.366 0.499**
after awareness (0.275) (0.231)
@@
Change in coefficient, -0.466 0.608**
after awareness (0.238) (0.240)
@@
Change in coefficient, -0.021** 0.023**
after awareness (0.010) (0.012)
@@
  • Notes: Dependent variable is defined as 40×fouls/minutes. Crew distance is the mean of the absolute distance in skin tone values between the player and each referee in the officiating crew. Each regression includes player-year fixed effects, game fixed effects, whether a player is a starter, whether a player is playing in her home arena, and her coach’s skin tone. The last two columns report the results of a pooled regression, including both a primary period and a subsequent period to measure the changes throughout our main sample years. Each observation is weighted by the number of minutes played. Standard errors are clustered at the player and game levels and appear in parentheses. ***p <0.01; **p <0.05; *p <0.1.
@@
Distribution of colorism by referee for the full sample period (2004-2014)

### v3

@@ 4.4  Individual referees
Since perceptions and consequences of colorism may differ by gender and skin tone of the referees [abramitzky2023, arceo2014, dixon2017, hill2002, thompson2001], it is interesting to explore individual, gender, or skin-tone-related referee bias or favoritism. Since we cannot precisely link individual referees to the exact fouls they call, we use the approach of [sandberg2018] by replacing the $\text{CrewDistance}_{ig}$ variable in equation (2) with an interaction term between a dummy for a specific referee and the absolute skin tone distance of each referee to a specific player, replacing $\gamma_{g}$ to control for attendance and team performance up until the specific game. We estimate these models separately for each referee to obtain coefficients indicating how much, on average, a referee deviates from the other referees. Overall, we do not find a clear pattern between the referees’ tone color and the players’ tone color and the foul calls. Likewise, no clear pattern emerges when looking at foul calling between female (triangles) and male (circles) referees (see Appendix Figures C.1, C.2 and C.3). [^24] Likewise, including a set of interaction terms regarding referee crew gender when estimating equation (2), we do not find evidence in our setting to suggest that referee gender influences the expression of colorism (results are available upon the request).
@@ 5  Discussion
In contrast to the results for the NBA by [price2010], we do not find evidence of significant and systematic racial discrimination in the WNBA before the widespread media coverage that was provoked by the publication of the [price2010] results. Remarkably, however, we observe that immediately following the increased public awareness (2007-2010), a player earns fewer fouls when facing more referees from the opposite race. While the estimates are not significant at conventional levels when using the FairFace algorithm, they are precise enough to become significant in the specifications using race categorizations from human raters. Moreover, across specifications, we find a clear and consistent pattern supporting this finding when using our continuous variable of skin tone.
Overall, this highlights a critical distinction between race and skin tone. In the binary classification by human raters and FairFace, race and skin tone are inherently entangled, and each evaluation relies on skin tone but also cues such as names and/or facial features to make racial assessments. In contrast, the continuous measure isolates skin tone. When doing so, significant results emerge, supporting the notion that skin tone is an important part of the mechanism and may warrant a reexamination of past empirical studies on racial discrimination.
The results in the NBA for the period immediately after the extensive media coverage are the same as in the WNBA, that is, players receive significantly fewer foul calls from the crew that has more opposite race referees. However, while the direction of a potential “media attention”-effect in the NBA and the WNBA is the same, the crucial difference between both settings is that before media exposure, there was evidence of racial discrimination in the NBA, while there was no such evidence in the WNBA. These ex-ante differences translate into zero racial bias in the NBA in the immediate post-awareness period, but overcorrection in the WNBA. [^25] The difference in results between the FairFace algorithm and human raters raises important questions about the broader implications of using state-of-the-art algorithms to make judgments traditionally reserved for humans. While we believe the analytical benefits, particularly in terms of scalability, replicability, and objectivity, far outweigh the potential drawbacks in this context, we acknowledge that these technologies introduce new ethical and practical considerations. Although a full discussion is beyond the scope of this paper, we believe there is a pressing need for ongoing dialogue about how best to integrate such advancements responsibly and effectively into empirical research.
@@ 5.1  Potential mechanisms
Our analyses allow us to probe deeper into the mechanisms of overcorrection. First, using the continuous variable of the skin tone, we find evidence that this overcorrection is driven by out-group favoritism rather than by in-group discrimination. This is in line with [feld2016], who investigate how student identity influences grader favoritism or discrimination and find a general pattern of favoritism rather than discrimination. Second, it is important to understand whether this out-group favoritism is intentional or not. For that, we dive into literature on the economics of discrimination, whose three main models include taste-based discrimination [becker1957], statistical discrimination [arrow1998, phelps1972], and implicit discrimination [bertrand2005].
The first type of model, taste-based discrimination, assumes that individuals have certain preferences against a certain group of individuals. However, this type is highly unlikely in a competitive market of sports referees [dohmen2008]. [price2010] also cast doubts about this mechanism in the NBA. In addition, our results regarding individual referees in the overcorrection period show that there are only very few referees with a significant bias, supporting the notion that taste-based discrimination is the unlikely candidate to explain our results. The second type of model, statistical discrimination, assumes asymmetric information where race is taken as a signal. As such, in the context of foul calls in the WNBA, where there is no problem of asymmetric information, this mechanism can be ruled out as well.
Consistent with other sports-related studies on referee bias, implicit discrimination is the most likely explanation for overcorrection in our case [price2010, gallo2013, ppw2018, faltings2023]. This alignment is supported by the fact that referee decisions in professional sports often involve high cognitive load (e.g., time pressure) and a degree of ambiguity, primary environmental preconditions for implicit bias [bertrand2005]. Given that we find evidence of out-group favoritism, we adjust the term from implicit discrimination to implicit favoritism. Such favoritism is unintentional and outside of the individual’s awareness. It may arise due to the unconscious association of a social group with different attributes and social norms. Social norms are especially important in the WNBA because of its strong emphasis on DEI. This commitment is apparent in the league’s programs, communications, and general atmosphere and thus may affect unconscious decisions that are taken in the type of split-second, high-pressure situations that appear in basketball games. [^26] It is worth mentioning a former player, Candice Wiggins, who claimed that the environment around the league was one of the reasons she chose to retire early. She mentioned feeling like an outsider in the WNBA as a heterosexual woman, noting that the league’s strong culture made it difficult for those who didn’t conform [wiggins2017].
@@ 5.2  Limitations
Despite the strengths of our approach, some limitations merit consideration. First, although we use a continuous, perceptually grounded measure of skin tone and validate racial classifications using both human raters and the FairFace algorithm, these methods depend on visual cues and are subject to noise from image quality, lighting, and subjective interpretation. While steps like HSV filtering and k-means clustering mitigate these issues, they cannot fully eliminate them. Second, our analysis relies on referee crew composition rather than individual-level foul call attribution. Thus, we cannot precisely link individual referees to the exact fouls they call, which would be ideal for isolating personal-level effects. Similarly, it would be informative to examine whether the observed favoritism arises because referees fail to call legitimate fouls or because they avoid making incorrect calls. Distinguishing between these possibilities would help determine whether overcorrection represents a genuine behavioral distortion or a more cautious, accuracy-driven adjustment in officiating, which itself warrants important future research to disentangle.
@@ 6  Conclusion
While increased awareness is supposed to reduce racial bias in organizations that have experienced racial discrimination, this paper aimed to investigate the effect of increased awareness in non-biased organizations regarding race and skin tone. For that, we used widespread media coverage of the racial bias of referees in the NBA to test whether it had an effect on the decision-making of referees in the WNBA, an organization that is known for its strong position on equal rights and fight against discrimination. Using state-of-the-art machine learning algorithms to define race and skin tone, our empirical analyses found that, as expected, WNBA referees were not systematically racially biased in their foul calls’ decisions before the media exposure of the NBA results. However, an increased awareness caused overcorrection, according to which players received fewer foul calls when facing more referees from the opposite race.
Despite the peculiar nature of the WNBA, our results may still have insight into the broader implementation of DEI initiatives. While we do not take a normative stance on such programs, rather, we emphasize that their impact depends heavily on context. In settings where baseline levels of bias are already low, well-intentioned interventions may unintentionally disrupt the already unbiased behaviors and create pressure to overcorrect. Moreover, while a significant body of economic research has explored out-group discrimination, our results call for a more comprehensive understanding of environments that may create overcorrection. Controlled experiments could test the generalizability of this phenomenon across other settings, while theoretical work could elucidate mechanisms linking group identity, salience, and behavior. Future research should aim to integrate these insights into models of social preferences, enriching our understanding of identity-driven decision-making in both individual and institutional contexts.
@@
  Table C.1 Racial bias among WNBA referees: Human-FairFace agreement .22
  Table C.2 Racial bias among WNBA referees: Team race control using FairFace .22
  Table C.3 Racial bias among WNBA referees: Team race control using human raters .23
  Table C.4 Racial bias among WNBA referees: FairFace classification balanced years .23
  Table C.5 Racial bias among WNBA referees: Human raters balanced years .23
  Table C.6 Colorism among WNBA referees: Balanced years .24
  Table C.7 Colorism among WNBA referees: Team skin-tone control .24
  Figure C.1 Distribution of colorism by referee (2004–06) .25
  Figure C.2 Distribution of colorism by referee (2007–10) .26
  Figure C.3 Distribution of colorism by referee (2011–14) .26
@@ Appendix B Methodology
As mentioned in the main text, our third approach follows the “Image-to-Data Pipeline” in [adukia2023], dividing the process into two main parts, i.e., face segmentation and skin tone classification. The detailed process is described here:
@@ B.1 Face segmentation
B.1.2 Isolating skin components [^27] We do, however, omit the Conditional Random Field (CRF), which is used by [adukia2023] to refine the skin mask by applying a CRF module, which predicts the labels of neighboring pixels (i.e., whether they are predicted to be skin or not skin) to produce a more fine-grained segmentation result. In our case, it is unnecessary since all our photos are human, and we will have additional skin mask refinement later (see feature processing). The added complexity a CRF module would bring is not a value add in our case scenario.: We use the Fully Connected Convolutional Neural Network (FC-CNN) method to extract a skin mask from faces. This involves applying a fully-convolutional neural network (FCN), a type of convolutional neural network (CNN) where the last fully connected layer is substituted with a convolutional layer that can capture the locations of the predicted labels. This enables the prediction of peripheral landmarks, such as the edges of the facial skin area, eyes, nose, and mouth.
@@
B.1.4 Creating face masks: Utilizing the predicted landmarks from the FC-CNN, we isolate the targeted facial region by extracting a skin “mask” from these predicted landmarks using the built-in Convex Hull method from SciPy’s Python library. The Convex Hull method generates the hull, and the skimage’s draw polygon function creates a binary mask isolating the face region.
@@ B.2.1 Skin Extraction
B.2.1.6 Conversion to BGR Color Space: The extracted skin regions were finally converted back to the BGR color space for analysis.
@@ Appendix C Results
Pre-awareness Post-awareness Change in coefficient
Black $\times$ 0.236 -0.264* 0.334* 0.237 -0.262*
fraction non-Black referees (0.204) (0.141) (0.174) (0.204) (0.141)
Post Black $\times$ -0.499* 0.593***
fraction non-Black referees (0.257) (0.211)
N 11,425 16,508 12,861 27,933 29,369
Sample mean 4.38 4.52 4.10 4.46 4.34
  • Notes: Estimated for a sub-sample restricted to cases where FairFace and human raters agree on racial classification. The dependent variable is defined as 40×fouls/minutes. Each regression includes player-year fixed effects, game fixed effects, whether a player is a starter, whether a player is playing in her home arena, and the skin color of her coach. The last two columns report the results of a pooled regression, including both a relevant primary period and a subsequent period to measure the change in our variable of interest between the two periods. Each observation is weighted by the number of minutes played. Standard errors are clustered at the player and game levels and appear in parentheses.
    ***p <0.01; **p <0.05; *p <0.1.
Table C.2:
Racial bias among WNBA referees using FairFace when controlling for team race
Pre-awareness Post-awareness Change in coefficient
Black $\times$ 0.077 -0.182 0.212 0.076 -0.181
fraction non-Black referees (0.206) (0.144) (0.165) (0.206) (0.144)
Post Black $\times$ -0.258 0.392*
fraction non-Black referees (0.262) (0.206)
N 12,893 17,266 15,705 30,159 32,971
  • Notes: Dependent variable is defined as 40×fouls/minutes. Each regression includes player-year fixed effects, game fixed effects, an indicator for whether the player is a starter, an indicator for whether the player is playing in her home arena, and the skin color of her coach. Also, a control for the player’s team’s racial composition (relative to the opponent) to account for differences in the on-court racial mix that could be correlated with officiating. The last two columns report pooled specifications that combine the primary period and subsequent period, allowing the coefficient on the variable of interest to differ across periods; the reported “post” interaction captures the change in the coefficient between the two periods. Each observation is weighted by minutes played. Standard errors are clustered at the player and game levels and reported in parentheses.
    ***p <0.01; **p <0.05; *p <0.1.
Table C.3:
Racial bias among WNBA referees using human raters when controlling for team race
Pre-awareness Post-awareness Change in coefficient
Black $\times$ 0.105 -0.258** 0.217 0.108 -0.257**
fraction non-Black referees (0.172) (0.129) (0.146) (0.172) (0.129)
Post Black $\times$ -0.364* 0.472**
fraction non-Black referees (0.220) (0.185)
N 12,893 17,266 15,705 30,159 32,971
  • Notes: See comments for Table C.2. ***p <0.01; **p <0.05; *p <0.1.
Table C.4:
@@
Post Black $\times$ -0.366 0.499**
fraction non-Black referees (0.275) (0.231)
@@
  • Notes: Estimated for even three-year intervals. The dependent variable is defined as 40×fouls/minutes. Crew distance is the mean of the absolute distance in skin tone values between the player and each referee in the officiating crew. Each regression includes player-year fixed effects, game fixed effects, whether a player is a starter, whether a player is playing in her home arena, and her coach’s skin tone. The last two columns report the results of a pooled regression, including both a primary period and a subsequent period to measure the changes throughout our main sample years. Each observation is weighted by the number of minutes played. Standard errors are clustered at the player and game levels and appear in parentheses.
    ***p <0.01; **p <0.05; *p <0.1.
@@
Post Black $\times$ -0.466 0.608**
fraction non-Black referees (0.238) (0.240)
@@
Post crew distance -0.021** 0.023**
(0.010) (0.012)
@@
  • Notes: See comments for Table C.4. ***p <0.01; **p <0.05; *p <0.1.
Table C.7:
Colorism among WNBA referees when controlling for team skin-tone
Pre-awareness Post-awareness Change in coefficient
Crew distance -0.000 -0.020∗∗∗ 0.006 -0.000 -0.020∗∗∗
(0.006) (0.007) (0.007) (0.006) (0.007)
Post crew distance -0.020∗∗ 0.026∗∗∗
(0.009) (0.010)
N 30,159 30,159 32,971 30,159 32,971
  • Notes: See comments for Table C.2. ***p <0.01; **p <0.05; *p <0.1.